Verify that the equation 2(y+z)dx - (x+z)dy+ (2y-x+z)dz = 0 is integrable and find its primitive .
1
Expert's answer
2019-08-22T15:54:13-0400

We have


"2(y+z)dx-(x+z)dy+(2y-x+z)dz=0"

General form of the Pfafffian equation is


"Pdx+Qdy+Rdz=0"

The integrability condition for the Pfaffian equation is


"(curlF, F)=0"

where "F=(P, Q, R)," or


"P({\eth Q \over \eth z}-{\eth R \over \eth y})+Q({\eth R \over \eth x}-{\eth P \over \eth z})+R({\eth P \over \eth y}-{\eth Q \over \eth x})=0"

Verify


"P=2(y+z), Q=-(x+z), R=(2y-x+z)""2(y+z)(-1-2)-(x+z)(-1-2)+(2y-x+z)(2+1)=""=-6y-6z+3x+3z+6y-3x+3z=0"

The integrability condition for this equation hold.

If the Pfaffian equation is multiplied by a certain function "\mu(x, y,z)" then one can obtain in the

left-hand side the total differential.

If we treat "z" as a constant, then the given equation reduces to


"2(y+z)dx-(x+z)dy=0""{2 \over x+z}dx-{1 \over y+z}dy=0"

which has a solution


"U(x, y, z)={(x+z)^2 \over y+z}=C_1"

The integrating factor "\mu(x,y,z)" is given by


"\mu={1 \over P}\cdot {\eth U \over \eth x}={1 \over 2(y+z)}\cdot {2(x+z) \over y+z}={x+z \over (y+z)^2}"

We have


"K=\mu R-{\eth U \over \eth z}=""={x+z \over (y+z)^2}(2y-x+z)-{2(x+z)(y+z)-(x+z)^2 \over (y+z)^2}=""={x+z \over (y+z)^2}(2y-x+z-2y-2z+x+z)=0"

Thus we have the equation


"dU=0"

which has a solution "U=c"

Thus the solution of the given equation


"(x+z)^2=c(y+z)"

where "c" is a constant.